# Extracting Sentence Embeddings from Pretrained Transformer Models

### v1

@@ 2.2.3 Retrofitting
This word vector space specialization method is used to incorporate semantic knowledge from external resources into word embeddings. The authors of [146] were the first to use the term “retrofitting” for their post-processing step of vector space word representations. Their idea was to incorporate rich relational information from semantic lexicons to word vectors, which are trained in a data-driven fashion using plain texts. The authors showed that the new vectors are similar in both their purely distributional representations and also in related word types. In addition, they showed improvements in several benchmark tasks.
Retrofitting can be useful for fitting the desired domain that is different from the original one used to pre-train word vectors. An example is [147] where the authors retrofitted to linkage information in biomedical taxonomy. Another related field is the alignment of cross-lingual word embeddings. The authors of [148] used retrofitting to align the vectors of the source and target languages in the dictionaries. In this way, translation pairs are pulled closer while minimizing deviation from the original embeddings and preserving the existing representation. In all of these cases, the retrofitting is based on external relational data.
According to the authors of [149], the weakness of retrofitting is that only the vectors of words present in the external constraints (resources) are modified. Therefore, they proposed an explicit retrofitting model in which external knowledge relations are turned into supervised training examples. Here, the distance between synonyms is supposed to be as low as possible, between antonyms as high as possible, while between remaining words not present in the external knowledge base it should remain the same as in the original representation space. This way, on top of word vectors, a deep feedforward neural network is trained to retrofit all word embeddings.
@@ 2.2.4 Other methods
Different post-processing can help surface the different information residing in the embeddings. The authors of [135] proposed a simple unsupervised singular value decomposition to reassign the feature weights. By changing the similarity order of their transformation, they tailored word embeddings in the semantics/syntax (tasks focusing whether on sing – chant or sing – singing) and similarity/relatedness (tasks solving car – automobile or car – road relationships) axes. Other authors of [152] used spectral filters to dissect BERT representations at different temporal scales. They showed that a low-pass filter yields the highest probing accuracy in topic classification, a high-pass filter in speech tagging, and a middle-pass filter is best in dialog act speech tasks.
@@ 2.2.5 Similarity measures
Some most basic tasks for sentences require measuring the distances between the corresponding embeddings. Usually, cosine similarity is employed. However, if the continuous representation space is curved or a text sequence is treated as a set of words, different metrics or representation space manipulations can be beneficial. The authors of [153] proposed Mutual Information (MI), well known in information theory and statistics, as a candidate for a similarity measure. They managed to successfully estimate MI for continuous random variables by the use of Kraskov–Stögbauer–Grassberger (KSG) estimator [154] which is based on elementary nearest-neighbor statistics. The other authors of [155] advised comparing sentence representations consisting of word vectors averaged by rank correlation, such as Kendall $\tau$. They argued that such measures enable mean pooled representations to rival modern deep ones, used with cosine similarity. The authors of [156] proposed treating sentences as fuzzy sets of words and showed good performance with the specially adapted DynaMax similarity measure. They showed that word vectors alone are sufficient to achieve excellent performance on semantic textual similarity tasks when sentence representations and similarity measures are derived using ideas from fuzzy set theory. Other authors of [157] suggest processing each sentence with respect to its found $k$-nearest neighbors. Then they find an optimal Euclidean subspace of the sentence manifold where cosine similarity would work the best. The authors of [158] also project the sentences onto a fixed-dimensional manifold with the objective of preserving local neighborhoods in the original space. All these mentioned works employ various strategies to enhance the comparison of document representations, tailoring their approaches to the given task at hand.
@@
Many similarity measures between sets employing earth mover’s distance were proposed. The authors of [159] were the first to frame the similarity of documents as a transportation problem. The idea is that the distances between similar but different words in two documents should be small. Their Word Mover’s Distance (WMD) is the cost of transporting a set of word vectors in an embedding space. This approach was effective, as it managed to exploit similarities between different word2vec word vectors, such as the relation of analogies. However, the main deficiency of WMD was that the distances were expensive to compute. Moreover, an output, the single number of a distance between two given documents, can be only combined with KNN or K-means, while usually applications require a whole feature vector. The authors of a subsequent work [160] managed to derive vectors for documents using WMD. They constructed a positive-definite word mover’s kernel using a feature map given by the WMD to random documents and then derived document embedding via a Random Features approximation of the kernel. Other authors of [70] proposed Word Rotator’s Distance (WRD). It is designed in a way that the norm and angle of word vectors correspond to the probability mass and transportation cost in the earth mover’s distance, respectively. Therefore, the norm of vectors, which is associated with the vector’s significance, does not interfere with the calculation of transportation cost, as in WMD. Finally, the authors of [161] attempted to include structural information that is absent in the WMD’s distance estimation between two sets of words. They represented the sentence vector as a weighted average of substructure vectors at a lower level in a recursive way, while transport plan at different level explain how the different substructures align.
2.3 Learning sentence representations directly
@@
The method resented in [162], and sometimes called doc2vec, is one of the first successful word2vec [16] adaptations to sequences of tokens. The authors presented two methods, of how document representation could be trained.
@@
In the Distributed Memory model of Paragraph Vectors (PV-DM), every paragraph is mapped to a unique vector which together with context words participates in the prediction of the next word. This way such a unique vector acts as a memory that stores the topic of the paragraph and bears its representation. The second is Distributed Bag Of Words version of Paragraph Vector (PV-DBOW) model. It leaves only a unique paragraph vector in the input and predicts words randomly sampled from that paragraph in the output. However, unlike the first method, it does not account for word ordering.
@@
In the original article [162], authors represented sentences as a concatenation of DM and DBOW vectors as they saw such a setup as more consistent. Later works [163, 164] also reported that individual model performance had only marginal differences, yet deviations were usually task-dependent.
@@
The authors of [165] proposed a modification to Paragraph Vectors and called it Doc2VecC. Differently from doc2vec, a document vector is derived not as a unique vector, but rather as an average of sampled constituent word embeddings. The other notable contribution is the added data-dependent regularization that favors informative or rare words while forcing the embeddings of common and non-discriminative ones to be close to zero. At the time the authors showed Doc2VecC to match state-of-the-art in multiple tasks.
To RNN- and transformer-based models
@@
Shallow doc2vec-like models soon met competition from more complicated Recurrent Neural Networks (RNN). In sequence-to-sequence [166] (like machine translation) setting such a model usually has two parts: encoder and decoder. The encoder, going through each token one by one, encodes the input sequence into a fixed hidden representation which is later used by the decoder to generate the target sequence (translation in NMT) token by token. Meanwhile, to generate sentence representation, only the decoder part is needed. Words of the sentence are sequentially fed as input to the RNN and the final hidden state is interpreted as a representation of the sentence.
@@
The benefit of RNN models is that the sequential consumption of tokens allows the final representation to account for the word order and process the arbitrary number of tokens. However, the effectiveness of this mechanism is questionable. It turns out that information has a hard time propagating all the way to the final hidden state. As compensation for that, a bidirectional [167] setting was used, which connects two hidden layers of opposite directions to the same output, simultaneously getting forward and backward information. The second solution was to use an attention mechanism [4], which during decoder token generation takes a weighted average of the encoder hidden states from all the input tokens. Yet it was only helpful for sequence-to-sequence tasks, like translation, as such a mechanism removed the last hidden state bottleneck, which was required to compress a long sequence of tokens into a single vector for a whole text representation.
@@
One of the first famous recurrent models for sentence representation is the SkipThought [168]. This model is trained to predict neighboring sentences from the source one. The center sentence is encoded by a bidirectional GRU (Gated Recurrent Units [169]) that concatenates the last hidden state of a forward GRU and the last hidden state of a backward GRU, and then decodes it into the two target sentences. This way the encoder is trained to map a sentence to a single representative vector. The authors showed that the model is robust and performed well on all tasks considered. The main drawbacks of this model are the month-long training, huge vector size of 4 800 (resulting from the concatenation of 2 400 vectors from two separate models), and support of only sentence-level representations, as well as the need for training text with a coherent inter-sentence narrative.
@@
As recurrent GRU or LSTM [170] architectures became the standard, the most recent advances in text representation were due to how data and training objectives were chosen. Here we have to mention a famous InferSent [12] model. The authors showed that the high-quality supervised data, although in low quantity, can give a huge increase in performance. They trained an encoder based on a bi-directional LSTM architecture with max pooling, on the Stanford Natural Language Inference (SNLI) [171] dataset. The resulting model outperformed SkipThought only after less than a day of training on a single GPU. Other authors of [172] modified the SkipThought model to QuickThoughts by framing the training task as a classification. Instead of generating the target sentence, the model encodes the likely candidates and chooses one of them. Authors of [173] leveraged several data sources with multiple training objectives. Their GenSen model training tasks included Neural Machine Translation, Constituency Parsing, Natural Language Inference, and SkipThought-like training. Finally, authors of [174] proposed even greater utilization of training data by constructing a discourse marker prediction task, predicting tokens such us tbecause, and, if, etc. Such framing of the task allowed the authors to mine vast amounts of text pairs together with the connecting markers. Their DisSent model achieved a similarly high performance to InferSent on a variety of transfer tasks. Overall, one can see that the way how model is trained, data quality, and quantity play a huge role in the resulting text representation performance.
@@
Many drawbacks of recurrent models were solved by the transformer architecture [2]. The authors decided to get rid of recurrence and leave only the attention mechanism [4] itself. The resulting model 1) is highly parallelizable; 2) only after the first layer each token has already attended to every other; and 3) is faster than RNN models when the sequence length is smaller than the representation dimensionality. These properties allowed us to train transformer models with staggering amounts of text much faster and resulted in state-of-the-art models on multiple tasks.
@@
Previous advances in sentence representation using RNNs were quickly applied to the new architecture. The authors of [175] presented a Universal sentence encoder model. It used only the encoder part of the original transformer model trained with multiple tasks: Skip-Thought-like training, SNLI, and conversational response prediction [176]. After the appearance of the pre-trained BERT model, the authors of [9] proposed SBERT: a pre-trained BERT model further fine-tuned with NLI data. It was a huge success, similar to the earlier InferSent model, yet this time both new architecture and general pre-training were the new factors for additional performance advancement. Other authors of [177] tried to further utilize the unsupervised data. Their PAUSE model learned sentence embeddings from a partially labeled dataset and showed that this way the same performance can be achieved with a smaller fraction of labeled NLI data. Here we see the same tendency as with the RNN models: new efforts to construct better fine-tuning tasks and better utilize existing data.
@@ 2.3.1 Contrastive learning approaches
Distance-based contrastive loss [178] is more attractive for learning sentence-level embeddings than more conventional error prediction losses. It allows for a simple construction of self-supervised learning using pairs of positive and negative examples. Embeddings of the first group samples are encouraged to be the same (by utilizing a loss on a distance function between the text vectors), while the negative ones are encouraged to be different.
In this way, the burden of expensive labeling can be relieved. Moreover, it can be used as an intermediate step between pre-training and fine-tuning to inexpensively align the model to the target task domain. These properties are especially useful for “data-hungry” transformer models.
@@
There are more sophisticated approaches to alter embeddings than a dropout randomly replacing weights with zero values. Specifically, adversarial attacks were shown to be successful. They aim to add worst-case perturbations to the input samples. The authors of [63] used Fast Gradient Value (FGV) [184] method, which unfortunately relies on supervised loss to compute adversarial perturbations. Other authors of [185] employed the Fast Gradient Sign Method (FGSM) [186]. Perturbation is obtained by applying the sign function to the derivative of the model with respect to the input by contrastive loss. The authors of [185] composed their best augmentation mix as a dropout with FGSM.
@@
Using synonym replacement, random swapping of two consequent words, random insertion and deletion, the so-called Easy Data Augmentation (EDA) the authors of [187] managed to reduce the required dataset size by half for the same performance in multiple classification tasks. In contrastive learning experiments for the biomedical relation extraction task the authors of [188] found that synonym replacement outperformed random swap and random deletion. Meanwhile, one of the first contrastive works for text [189] showed that back-translation [190] augmentation outperformed the four mentioned techniques. The authors took a pre-trained BERT [5] and pre-trained it further on the input text of GLUE [191] tasks in a contrastive fashion, to be later fine-tuned and evaluated in GLUE. The resulting CERT model achieved slight improvements in addition to the regular BERT. CLEAR [192] was also an attempt to improve the regular BERT and RoBERTa pre-training by jointly using MLM loss and constrastive loss. The authors found that the different combinations of different augmentation strategies (in particular, substitution, span-deletion, and reordering) favor different improvements for GLUE and various SentEval tasks. Recent work of [185] tried token-level augmentations such as typos, synonym replacement, paraphrasing, and back-translation but found that they underperformed feature-level ones.
As the popular BERT model is pre-trained with masked language modeling task, i.e., predicting the true masked token behind the special input [MASK] token, the same [MASK] can be also used to induce augmentations by randomly masking tokens. The mirror-BERT model of [180] used such random span masking for input augmentation, which together with feature augmentation showed gains over off-the-shelf models in both lexical and sentence level tasks, across different domains and different languages.
@@
A better distinction of negatives can also be made by token augmentations. The authors of [92] proposed Bidirectional Margin Loss (BML) to incorporate soft-negative samples that are generated using a simple rule-based method. According to its dependency syntax tree, the positive sentence is converted to its negation with correct syntax and clear semantics. The proposed setup saw improvements in the semantic textual similarity tasks from SentEval.
@@
Similar to the older idea of skip-thoughts [168], the placement of sentences in a text can be exploited. In such a setting, text segments near or overlapping each other in the long text should also be near each other in the text embedding space. Therefore, text parts near the same anchor in the text can be regarded as positive, while further away as negative pairs for contrastive learning. DeCLUTR [193] for each document in a mini-batch samples multiple anchor spans. The corresponding positives are further sampled for each anchor and can be partially overlapping, adjacent, or subsumed by the anchor. This is accomplished by sampling the anchor span length to be mostly longer than the positive spans. At the time, DeCLUTR showed improvements in SentEval [194] transfer tasks in an unsupervised setting. Location exploitation is a simple approach, but it has some drawbacks. As the authors used the dataset consisting of documents of at least 2 048 tokens in length, it clearly reveals that sentence placement methods have limited applications for short text domains. Moreover, random spans of DeCLUTR are subject to fragmentation in semantics.
@@
The augmented version of the sample can also be constructed using two encoders. Different weight initialization and training data order are the most common causes of fluctuations in neural network performance. Although general performance is of a comparable level, each model learned differently represents the different local minimum. Such a subtle difference can act as an augmentation. One of the first such setups was the CT model [99]. The simple objective required two models to retain similar representations for identical sentences while distinguishing their representations from different ones.
@@
The authors of [195] proposed to use multiple diverse positives instead of the usual two. The authors claim that this increases the probability of “at-least-one” effective positive during training. To implement diverse positives, the agreement of two similarity distributions of samples between two encoder models had to be maximized.
Another setup incorporating two networks can be used for Knowledge Distillation (KD). In [196], after the regular KD step with the original pre-training task, an additional step with a contrastive pre-training was added. That is, the pair of positives for a single text segment consisted of a representation from the teacher model and another from the student. After the third stage of fine-tuning on semantic textual similarity tasks [196] showed that the 110M model outperformed the one with 11B parameters.
@@
One of the first works to implement this idea was SimSCE [143]. The authors analyzed six candidate datasets. They found that the Natural Language Inference (NLI) datasets, SNLI [171] and MNLI [197], together performed the best. The combined dataset contains sentence pairs in the form of (premise, hypothesis, label), labeled as entailment, neutral, or contradiction, 314k for each class. SimSCE authors constructed positive pairs from entailment samples and negatives using contradiction ones. The resulting supervised approach reached state-of-the-art results at the time for SentEval tasks. The authors also tried to incorporate neutral pairs as less weighted negatives but did not observe improvements.
@@
ST5 [200] adopted the two-stage training strategy. In the first, 2B mined question-answering data from Community QA sites were used framing the question and answer into the positive pair. During the second stage, NLI data was used, similar to SimSCE [143]. Such a more data-rich training allowed the authors to outperform the previous approaches for the SentEval task. Furthermore, the use of larger models up to 11B parameters brought even further gains.

### v2

@@ 2.2.3 Retrofitting
This word vector space specialization method can incorporate semantic knowledge from external resources into word embeddings. The authors of [146] were the first to use the term “retrofitting” for their post-processing step of vector space word representations. Their idea was to incorporate rich relational information from semantic lexicons to word vectors, which are trained in a data-driven fashion using plain texts. The authors showed that the new vectors are similar in both their purely distributional representations and related word types. In addition, they showed improvements in several benchmark tasks.
Retrofitting can be useful for fitting a desired domain that is different from the original one used to pre-train the word vectors. An example is [147] where the authors retrofitted to linkage information in biomedical taxonomy. Another related field is the alignment of cross-lingual word embeddings. The authors of [148] used retrofitting to align the vectors of the source and target languages in the dictionaries. This way, translation pairs are pulled closer while minimizing deviation from the original embeddings and preserving the existing representation. In all these cases, the retrofitting is based on external relational data.
According to the authors of [149], the weakness of retrofitting is that only the vectors of words present in the external constraints (resources) are modified. Therefore, they proposed an explicit retrofitting model in which external knowledge relations are turned into supervised training examples. Here, the distance between synonyms is supposed to be as low as possible, and between antonyms as high as possible, while between remaining words not present in the external knowledge base, it should remain the same as in the original representation space. This way, on top of word vectors, a deep feedforward neural network is trained to retrofit all the word embeddings.
@@ 2.2.4 Other Methods
Different post-processing can help surface the different information residing in the embeddings. The authors of [135] proposed a simple unsupervised singular value decomposition to reassign the feature weights. By changing the similarity order of their transformation, they tailored word embeddings in the semantics/syntax (tasks focusing on sing–chant or sing–singing) and similarity/relatedness (tasks solving car–automobile or car–road relationships) axes. Other authors of [152] used spectral filters to dissect BERT representations at different temporal scales. They showed that a low-pass filter yields the highest probing accuracy in topic classification, a high-pass filter in speech tagging, and a middle-pass filter is best in dialog act speech tasks.
@@ 2.2.5 Similarity Measures
Some most basic tasks for sentences require measuring the distances between the corresponding embeddings. Usually, cosine similarity is employed. However, if the continuous representation space is curved or a text sequence is treated as a set of words, different metrics or representation space manipulations can be beneficial. The authors of [153] proposed Mutual Information (MI), well known in information theory and statistics, as a candidate for a similarity measure. They managed to successfully estimate MI for continuous random variables by the use of Kraskov–Stögbauer–Grassberger (KSG) estimator [154], which is based on elementary nearest-neighbor statistics. The other authors of [155] advised comparing sentence embeddings consisting of word vectors averaged by rank correlation, such as Kendall’s $\tau$. They argued that such measures enable mean pooled representations to rival modern deep ones, used with cosine similarity. The authors of [156] proposed treating sentences as fuzzy sets of words and showed good performance with the specially adapted DynaMax similarity measure. They showed that word vectors alone are sufficient to achieve excellent performance on semantic textual similarity tasks when sentence embeddings and similarity measures are derived using ideas from fuzzy set theory. Other authors of [157] suggest processing each sentence with respect to its found $k$-nearest neighbors. Then, they find an optimal Euclidean subspace of the sentence manifold where cosine similarity would work best. The authors of [158] also project the sentences onto a fixed-dimensional manifold with the objective of preserving local neighborhoods in the original space. All these works mentioned employ various strategies to enhance the comparison of document representations, tailoring their approaches to the given task.
@@
Many similarity measures between sets employing earth mover’s distance were proposed. The authors of [159] were the first to frame the similarity of documents as a transportation problem. The idea is that the distances between similar but different words in two documents should be small. Their Word Mover’s Distance (WMD) is the cost of transporting a set of word vectors in an embedding space. This approach was effective, as it managed to exploit similarities between different Word2Vec word vectors, such as the relation of analogies. However, the main deficiency of WMD was that the distances were expensive to compute. Moreover, an output, the single number of a distance between two given documents, can only be combined with $k$-nearest neighbors or $k$-means, while applications usually require a whole feature vector. The authors of a subsequent work [160] managed to derive vectors for documents using WMD. They constructed a positive-definite word mover’s kernel using a feature map given by the WMD to random documents and then derived document embedding via a Random Features approximation of the kernel. Other authors of [70] proposed Word Rotator’s Distance (WRD). It is designed so that the norm and angle of word vectors correspond to the probability mass and transportation cost in the earth mover’s distance, respectively. Therefore, the norm of vectors, which is associated with the vector’s significance, does not interfere with the calculation of transportation cost, as in WMD. Finally, the authors of [161] attempted to include structural information absent in the WMD’s distance estimation between two sets of words. They represented the sentence vector as a weighted average of substructure vectors at a lower level in a recursive way, while a transport plan at a different level explains how the different substructures align.
2.3 Learning Sentence Embeddings Directly
@@ 2.3.1 Paragraph Vectors
The method presented in [162], and sometimes called Doc2Vec, is one of the first successful Word2Vec [16] adaptations to sequences of tokens. The authors presented two methods of how document representation could be trained.
@@
In the Distributed Memory model of Paragraph Vectors (PV-DM), every paragraph is mapped to a unique vector, which, together with context words, participates in predicting the next word. This way, such a unique vector acts as a memory that stores the topic of the paragraph and bears its representation. The second is the Distributed Bag Of Words version of the Paragraph Vector (PV-DBOW) model. It leaves only a unique paragraph vector in the input and predicts words randomly sampled from that paragraph in the output. However, unlike the first method, it does not account for word ordering.
@@
In the original article [162], authors represented sentences as a concatenation of DM and DBOW vectors, as they saw such a setup as more consistent. Later works [163, 164] also reported that individual model performance had only marginal differences, yet deviations were usually task-dependent.
@@
The authors of [165] proposed a modification to Paragraph Vectors and called it Doc2VecC. Differently from Doc2Vec, a document vector is derived not as a unique vector, but as an average of sampled constituent word embeddings. The other notable contribution is the added data-dependent regularization that favors informative or rare words while forcing the embeddings of common and non-discriminative ones to be close to zero. At the time, the authors showed Doc2VecC to match state-of-the-art in multiple tasks.
2.3.2 To RNN- and Transformer-Based Models
@@
Shallow Doc2Vec-like models soon met competition from more complicated Recurrent Neural Networks (RNNs). In a sequence-to-sequence [166] (like machine translation) setting, such a model usually has two parts: encoder and decoder. The encoder, going through each token one by one, encodes the input sequence into a fixed hidden representation, which is later used by the decoder to generate the target sequence (translation in NMT) token by token. Meanwhile, only the decoder part is needed to generate sentence representation. Words of the sentence are sequentially fed as input to the RNN, and the final hidden state is interpreted as its representation.
@@
The benefit of RNN models is that the sequential consumption of tokens allows the final representation to account for the word order and process the arbitrary number of tokens. However, the effectiveness of this mechanism is questionable. It turns out that information has a hard time propagating all the way to the final hidden state. As compensation for that, a bidirectional [167] setting was used, which connects two hidden layers of opposite directions to the same output, simultaneously getting forward and backward information. The second solution was to use an attention mechanism [4], which, during decoder token generation, takes a weighted average of the encoder hidden states from all the input tokens. But it was only helpful for sequence-to-sequence tasks, like translation, as such a mechanism removed the last hidden state bottleneck, which was required to compress a long sequence of tokens into a single vector for a whole text representation.
@@
One of the first famous recurrent models for sentence representation is the SkipThought [168]. This model is trained to predict neighboring sentences from the source one. The center sentence is encoded by a bidirectional GRU (Gated Recurrent Unit [169]) that concatenates the last hidden state of a forward GRU and the last hidden state of a backward GRU, and then decodes it into the two target sentences. This way, the encoder is trained to map a sentence to a single representative vector. The authors showed that the model is robust and performed well on all tasks considered. The main drawbacks of this model are the month-long training, huge vector size of 4800 (resulting from the concatenation of 2400 vectors from two separate models), and support of only sentence-level embeddings, as well as the need for training text with a coherent inter-sentence narrative.
@@
As recurrent GRU or LSTM [170] architectures became the standard, the most recent advances in text representation were due to how data and training objectives were chosen. Here, we have to mention a famous InferSent [12] model. The authors showed that the high-quality supervised data, although in low quantity, can hugely increase the performance. They trained an encoder based on a bi-directional LSTM architecture with max pooling, on the Stanford Natural Language Inference (SNLI) [171] dataset. The resulting model outperformed SkipThought after less than a day of training on a single GPU. Other authors of [172] modified the SkipThought model to QuickThoughts by framing the training task as a classification. Instead of generating the target sentence, the model encodes the likely candidates and chooses one of them. Authors of [173] leveraged several data sources with multiple training objectives. Their GenSen model training tasks included Neural Machine Translation, Constituency Parsing, Natural Language Inference, and SkipThought-like training. Finally, authors of [174] proposed even greater utilization of training data by constructing a discourse marker prediction task, predicting tokens such us because, and, if, etc. Such framing of the task allowed the authors to mine vast amounts of text pairs together with the connecting markers. Their DisSent model performed similarly well to InferSent on various transfer tasks. Overall, one can see that how the model is trained and data quality and quantity play a huge role in the resulting text representation performance.
@@
Many drawbacks of the recurrent models were solved by the transformer architecture [2]. The authors decided to get rid of recurrence and leave only the attention mechanism [4] itself. The resulting model (1) is highly parallelizable; (2) only after the first layer, each token has already attended to every other; and (3) it is faster than RNN models when the sequence length is smaller than the representation dimensionality. These properties allowed us to train transformer models with staggering amounts of text much faster and they became state-of-the-art models on multiple tasks.
@@
Previous advances in sentence representation using RNNs were quickly applied to the new architecture. The authors of [175] presented a Universal sentence encoder model. It used only the encoder part of the original transformer model trained with multiple tasks: Skip-Thought-like training, SNLI, and conversational response prediction [176]. After the appearance of the pre-trained BERT model, the authors of [9] proposed SBERT: a pre-trained BERT model further fine-tuned with NLI data. It was a huge success, similar to the earlier InferSent model, yet this time both the new architecture and the general pre-training were the new factors for additional performance advancement. Other authors of [177] tried to further utilize the unsupervised data. Their PAUSE model learned sentence embeddings from a partially labeled dataset and showed that this way, the same performance can be achieved with a smaller fraction of labeled NLI data. Here, we see the same tendency as with the RNN models: new efforts to construct better fine-tuning tasks and better utilize existing data.
@@ 2.3.3 Contrastive Learning Approaches
Distance-based contrastive loss [178] is more attractive for learning sentence-level embeddings than the more conventional error prediction losses. It allows for a simple construction of self-supervised learning using pairs of positive and negative examples. Embeddings of the first group of samples are encouraged to be the same (by utilizing a loss on a distance function between the text vectors), while the negative ones are encouraged to be different.
In this way, the burden of expensive labeling can be relieved. Moreover, it can be used as an intermediate step between pre-training and fine-tuning to inexpensively align the model to the target task domain. These properties are especially useful for the “data-hungry” transformer models.
@@
There are more sophisticated approaches to alter embeddings than a dropout randomly replacing weights with zero values. Specifically, adversarial attacks were shown to be successful. They aim to add worst-case perturbations to the input samples. The authors of [63] used the Fast Gradient Value (FGV) [184] method, which unfortunately relies on supervised loss to compute adversarial perturbations. Other authors of [185] employed the Fast Gradient Sign Method (FGSM) [186]. Perturbation is obtained by applying the sign function to the derivative of the model with respect to the input by contrastive loss. The authors of [185] composed their best augmentation mix as a dropout with FGSM.
@@
Using synonym replacement, random swapping of two consequent words, random insertion and deletion, the so-called Easy Data Augmentation (EDA) the authors of [187] managed to reduce the required dataset size by half for the same performance in multiple classification tasks. In contrastive learning experiments for the biomedical relation extraction task, the authors of [188] found that synonym replacement outperformed random swap and random deletion. Meanwhile, one of the first contrastive works for text [189] showed that back-translation [190] augmentation outperformed the four mentioned techniques. The authors took a pre-trained BERT [5] and pre-trained it further on the input text of GLUE [191] tasks in a contrastive fashion, to be later fine-tuned and evaluated in GLUE. The resulting CERT model achieved slight improvements in addition to the regular BERT. CLEAR [192] was also an attempt to improve the regular BERT and RoBERTa pre-training by jointly using MLM loss and constrastive loss. The authors found that the different combinations of different augmentation strategies (in particular, substitution, span-deletion, and reordering) favor different improvements for GLUE and various SentEval tasks. A recent work [185] tried token-level augmentations such as typos, synonym replacement, paraphrasing, and back-translation but found that they underperformed feature-level ones.
As the popular BERT model is pre-trained with a masked language modeling task, i.e., predicting the true masked token behind the special input [MASK] token, the same [MASK] can also be used to induce augmentations by randomly masking tokens. The mirror-BERT model of [180] used such random span masking for input augmentation, which, together with feature augmentation, showed gains over off-the-shelf models in both lexical and sentence-level tasks, across different domains and different languages.
@@
A better distinction of negatives can also be obtained by token augmentations. The authors of [92] proposed Bidirectional Margin Loss (BML) to incorporate soft-negative samples that are generated using a simple rule-based method. According to its dependency syntax tree, the positive sentence is converted to its negation with correct syntax and clear semantics. The proposed setup saw improvements in the semantic textual similarity tasks from SentEval.
@@
Similar to the older idea of skip-thoughts [168], the placement of sentences in a text can be exploited. In such a setting, text segments near or overlapping each other in the long text should also be near each other in the sentence embedding space. Therefore, text parts near the same anchor in the text can be regarded as positive, while further away as negative pairs for contrastive learning. DeCLUTR [193] for each document in a mini-batch samples multiple anchor spans. The corresponding positives are further sampled for each anchor and can be partially overlapping, adjacent, or subsumed by the anchor. This is accomplished by sampling the anchor span length to be mostly longer than the positive spans. At the time, DeCLUTR showed improvements in SentEval [194] transfer tasks in an unsupervised setting. Location exploitation is a simple approach, but it has some drawbacks. As the authors used the dataset consisting of documents of at least 2048 tokens in length, it clearly reveals that sentence placement methods have limited applications for short text domains. Moreover, random spans of DeCLUTR are subject to fragmentation in semantics.
@@
The augmented version of the sample can also be constructed using two encoders. Different weight initialization and training data order are the most common causes of fluctuations in neural network performance. Although the performance is generally comparable, each model learned differently represents a different local minimum. Such a subtle difference can act as an augmentation. One of the first such setups was the CT model [99]. The simple objective required two models to retain similar representations for identical sentences while distinguishing their representations from different ones.
@@
The authors of [195] proposed using multiple diverse positives instead of the usual two. The authors claim that this increases the probability of “at-least-one” effective positive during training. To implement diverse positives, the agreement of two similarity distributions of samples between two encoder models had to be maximized.
Another setup incorporating two networks can be used for Knowledge Distillation (KD). In [196], after the regular KD step with the original pre-training task, an additional contrastive pre-training was added. The pair of positives for a single text segment consisted of a representation from the teacher model and another from the student. After the third stage of fine-tuning on semantic textual similarity tasks, [196] showed that the 110 M model outperformed the one with 11B parameters.
@@
One of the first works to implement this idea was SimSCE [143]. The authors analyzed six candidate datasets. They found that the Natural Language Inference (NLI) datasets, SNLI [171] and MNLI [197], together performed the best. The combined dataset contains sentence pairs in the form of (premise, hypothesis, label), labeled as entailment, neutral, or contradiction, 314k for each class. The SimSCE authors constructed positive pairs from entailment samples and negatives using contradiction ones. The resulting supervised approach achieved state-of-the-art results at the time for SentEval tasks. The authors also tried to incorporate neutral pairs as less weighted negatives but did not observe improvements.
@@
ST5 [200] adopted the two-stage training strategy. In the first, 2B mined question-answering data from Community QA sites were used, framing the question and answer into the positive pair. During the second stage, NLI data were used, similar to SimSCE [143]. Such a more data-rich training allowed the authors to outperform the previous approaches for the SentEval task. Furthermore, using larger models with up to 11B parameters brought even further gains.